OpenStudy (anonymous):

Choose the correct simplification of the expression (−11x3y12z3)0. zero 6x3y12z3 −11 1 12. [7.06] Choose the correct simplification of the expression −6x2(4x − 7x2 − 2). 42x4 + 24x3 − 12x2 −42x4 + 24x3 − 12x2 −13x4 − 2x3 − 8x2 42x4 − 24x3 + 12x2

Nnesha (nnesha):

(jo2398)^0 =1 anything to the 0 power = just 1

Nnesha (nnesha):

x^0 = ? u^0 = ?? (7780a^3 b^7)^0 = ?